OpenStudy (azureilai):

help in trig identities please! cos4xcosx-sin4xsinx=0 My calculus friend came up with the answer of (pi/2)-x=4x, resulting in x=pi/10. I graphed it, and it was an answer within the interval of [0,2pi] But he didn't explain his work on how he arrived at that answer. Can someone help please?

OpenStudy (accessdenied):

Do you know of a trig identity that has this form? __ = cos a cos b - sin a sin b

OpenStudy (azureilai):

nope, which identity is that?

OpenStudy (azureilai):

When I was trying to solve this, I tried using the double angle formulas, but I couldn't get it to work.

OpenStudy (anonymous):

\[\cos \left( a+b \right)=\cos a \cos b-\sin a \sin b\]

OpenStudy (azureilai):

so would that be applying it to \[\cos 4x=\cos2\cos2-\sin2\sin2\]?

OpenStudy (anonymous):

\[\cos 4x \cos x-\sin 4x \sin x=\cos \left( 4x+x \right)=\cos 5x=0\]

OpenStudy (azureilai):

Oh ok. I see how thats applied now. Thanks, it makes the problem alot more solvable once sin is out.

OpenStudy (anonymous):

\[\cos 5x=0=\cos \left(2 n+1 \right)\frac{ \pi }{2 },5x=\left( 2n+1 \right)\frac{ \pi }{ 2 }\] \[x=\left( 2n+1 \right)\frac{ \pi }{ 10 }\],n=0,1,2,...,9

OpenStudy (azureilai):

Ok, I got it know thanks alot for your help.

OpenStudy (anonymous):

yw